Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on October 11th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • DC report
  • Magnet status
    • Daresbury: Investigating SSD SD-01 sensor on HTS leads
    • Daresbury: Investigate SSU M1 trip
    • Daresbury: Commission SS water solenoids for diode absorbers
    • Daresbury: Calibrate dipoles (conventional magnets)
  • LiH absorber installation status
  • Decay solenoid status
    • Daresbury: Replace battery and fit diodes to charging circuit
    • Daresbury: Recalibrate PSU
  • Water status
    • Daresbury: Check control & monitoring
  • Wednesday maintenance jobs 12th October
  • Shutdown activities : 28 October - 15 November
    • Mop-up of summer list and other issues?
    • SSD cryocooler system refurb (Nov 9)
    • Possible test of new Linde control PC
    • Conventional magnets: Earth leakage tests at magnets (inc. synch and DSA)

Operations

  • MOM report
    • When do shifts begin on CHEESE : Every "day" slot on CHEESE starts at 08:00 and continues until 08:00 the next calendar day.
  • Detectors status (1 minute/1 slide)
    • Melissa - Tracker
    • Mariyan - KL
    • Scott - TOF
    • Francois - EMR
    • Lucien - Ckov
    • Rogers - Physics
  • BLOC for last shift block (Oct 23 - Oct 28)
  • Magnet monitoring shifts from Oct 15 to end of running

Minutes

Present: S. Boyd, M. Tucker, Y Song, M. Popovic, A. Young, P. Franchini, K. Long, D. Sanders, P. Kyberd,
C. Rogers, V. Blackmore, V. Bayliss, H. Nebrensky, C Whyte, J Boehm, P Hanlet

  • DC report
    It's busy. Hall workplan exists.
  • Magnet status
    • Daresbury: Investigating SSD SD-01 sensor on HTS leads
      • Done. Traced back to the V3 or V1 feedthrough. Looks like it's an internal problem. No consquences due to sensor redundancy.
        Colin : any reason for this change? Pierrick : one of them was flaky in the past but can't remember which one. It may be a fluctuating bad sensor but history doesn't exist.
    • Daresbury: Investigate SSU M1 trip
      • M1 coil of SSU. Reproducible as of this weekend. Appeared with 6.1.4 config and then ramped to another config (successfully) and then it tripped. Everything ramped down and M1 coil only ramped. It tripped at 28A. Then 11 A. Turned of PS on the weekend and tried to reproduce the error. It was not reproduced. Daresbury swapped out controller and power supply but won't know until we can ramp if this has worked.
        Josef : Daresbury should take PS with them to Daresbury and check it out. Colin : this may not be possible since it may be an interlock problem. Pierrick : We do have spare PS's. Colin : in which case we should make Daresbury look at it. Ken : we should just put it in a cabinet and mark it suspicious until we know whether the PS or controller was at fault.
    • Daresbury: Commission SS water solenoids for diode absorbers
      • This is the emergency water source changeover. Not done.
    • Daresbury: Calibrate dipoles (conventional magnets)
      • Hasn't been done this time since they are working mainly on the solenoids. Henry things the dipole job can be delegated to
        Richard Hale.
  • LiH absorber installation status
    • Very nearly ready to move. Could be offline tonight.
    • Installation estimate : "The job will take up to a week. Hoping to be largely done by end of Friday".
    • Do we need to survey the system after installation? Yes. Victoria will arrange a survey.
    • Pierrick : when the hoses are configured could it be such that they don't cover the readout and control boxes for the heater tape on the SS's. Colin : we'll do our best.
  • Decay solenoid status
    • Problem with the level sensor which was reading all over the place. Miraculously worked.
    • Is the DS ready to run? Yes, but if the level sensor goes wrong the current will ramp down.
    • Ken : can we run it at 200 A overnight and soak test it.
    • This will need to be ramped down early in the morning. Vicky will turn it off before 9 am.
  • Daresbury: Replace battery and fit diodes to charging circuit.
    • Will check
  • Daresbury: Recalibrate PSU
    • In hand. Should be done today.
  • Water status
    • Daresbury: Check control & monitoring
  • Colin : guy who was going to work on the manager that controls the two coolers on the roof didn't turn up. However,
    don't know what is going on yet. Josef : one cause could be too high a glycol concentration. Colin : yes that could
    be an issue.
    Henry : can we please monitor the temperatures on shift.
    After extended discussion : limit will be lowered in alarm handler and instructions for shifters made clear.
    Conventional magnet water temperature in ALH on Thursday.
    Henry : Highest acceptable momentum setting was 240 MeV. Check with Henry if we want to go higher. Run plan needs to
    be managed so "unconventional" momentum settings are considered whilst Henry is available (9-6 on weekdays)
    Colin : we need to get the conventional magnet temperatures on the GUIs. Pierrick is doing this at the moment.
  • Wednesday maintenance jobs 12th October
    • Is is possible that two TOF2 channels have been swapped in the maps. Need to check this with Durga.
  • Shutdown activities : 28 October - 15 November
    • Mop-up of summer list and other issues?
    • SSD cryocooler system refurb (Nov 9)
    • Possible test of new Linde control PC
    • Conventional magnets: Earth leakage tests at magnets (inc. synch and DSA)

Operations

  • MOM report
    • When do shifts begin on CHEESE : Every "day" slot on CHEESE starts at 08:00 and continues until 08:00 the next calendar day.
  • BLOC for last shift block (Oct 23 - Oct 28)
  • Magnet monitoring shifts from Oct 15 to end of running

InstructionsAndPics.pdf - Helium Instructions and Pics (38.8 MB) George, Melissa, 11 October 2016 00:08

OpsMeet11_10_2016.pdf - Tracker Data Reconstruction update (215 KB) George, Melissa, 11 October 2016 00:36

Run8376imageData.pdf - Run 8376, SSU+SSD+FC (965 KB) Rajaram, Durga, 11 October 2016 06:50

Run8407imageData.pdf - Run 8407, Straight tracks (788 KB) Rajaram, Durga, 11 October 2016 06:51